inverse of y = √x+1 +√x-1
please can someone help with this ?

Grade:12th pass

2 Answers

Arun
5 years ago
Dear student
 
Just replace x by y and y by x.
 
hence inverse will be
 
x = √y+1 +√y-1
 
Hope it helps
 
Regards
Arun ( askIITians forum expert)
 
GUDISA PRASUNA
5 years ago
dear friend
f(x)= √x+1 +√x-1 
y=2√x
y/2=(√x)
(y/2)^2=x
then in the last equation replace x with y and y with x
f’(x)=(x/2)^2